OpenStudy (anonymous):

A box contains 12oz of cereal. Each serving of cereal is 3/4 oz. How many servings are contained in the box?

OpenStudy (stamp):

1 serving = 3/4 oz 1 box = 12 oz Observe. (1 serving = 3/4 oz) * 4 leads to 4 servings = 3 oz so (4 servings = 3 oz) * 4 is 16 servings = 12 oz finally 16 servings = 12 oz 1 box = 12 oz equate the two to obtain 1 box = 16 servings

OpenStudy (anonymous):

16 12 / .75 = 16